OpenStudy (anonymous):

Solve (1/6x^2) = ( 1/3x^2) - (1/x)

OpenStudy (jhannybean):

\[\large \frac{1}{6x^2}= \frac{1}{3x^2}-\frac{1}{x}\] make a common denominatoron the RHS

OpenStudy (dan815):